A transverse wave on a taut string is modeled with the wave function y(x, t) = Asin(kx-wt) (0.2m)sin(6.28m-1 x – 1.57s-1 t) Find the: (a) amplitude, (b) wavelength, (c) period, (d) speed of the wave.
1
Expert's answer
2021-10-02T17:34:34-0400
"A=0.2\ m\\\lambda=\frac{2\pi}{6.28}=1\ m\\ T=\frac{2\pi}{1.57}=4\ s\\v=\frac{1}{4}=0.25\frac{m}{s}"